OpenStudy (anonymous):

What is the value of the expression shown? 10 – 16 ÷ 2 A -3 B 2 C 3 D 18

OpenStudy (baby_bear69):

Always do multiplication or division first. so then it would be 10-8=?

OpenStudy (anonymous):

16/2=8 10-8=2 2=B

OpenStudy (sohailiftikhar):

yes it's B

pooja195 (pooja195):

Well you should've used PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction first comes division 16/2=8 then subtraction 10-8=2

OpenStudy (anonymous):

2